OpenStudy (owlet):

Find the intersection of \(x^3\) and \(3^x\) How to do it?

OpenStudy (zzr0ck3r):

Hint for one solution: \(a^a=a^a\). The other solution gets a little tricky. Google "analytic continuation of the product log function"

OpenStudy (owlet):

so x=3? one point of intersection would be (3, 27) ?

OpenStudy (zzr0ck3r):

Are you asking me if 3^3 = 27?

OpenStudy (owlet):

no, if one of the intersection points of these two function would be (3, 27)

OpenStudy (owlet):

how about the other point?

OpenStudy (zzr0ck3r):

put in \(x=3\) on both functions, and then see what you get out...

OpenStudy (zzr0ck3r):

As far as the other point, I refer you to my first comment.

OpenStudy (zzr0ck3r):

The solution is transcendental, so like \(\pi\) there is no good way to explain it with numbers :)

OpenStudy (owlet):

oh got it. thanks.. no point of solving this thing because its too complicatedlol but I understand the first solution though :) tysm

OpenStudy (zzr0ck3r):

yeah, it is crazy. There actually might be infinite solutions, but only one is algebraic, meaning we can express it with numbers and symbols...